# Past Movements-Guided Motion Representation Learning for Human Motion Prediction

### v1

@@ 3.4 Fine-tuning for Motion Prediction
$\mathbf{H}^{F}$ $=\mathcal{F}_{JE}(\mathcal{F}_{FMP}(\mathbf{H}_{0}^{F},\mathbf{H}^{P}))$ (6)
$\mathbf{X}^{F}$ $=\mathcal{F}_{PH}(\mathbf{H}^{F})$ (7)
where $\mathbf{H}_{0}^{F}$ is the initial state initialized with the value of 0. $\mathcal{F}_{JE}$ represents the joint embedder.
@@ 3.5 Loss Function
Loss Function of Pretrain Stage
Pretrain stage performs past sequence reconstruction and future sequence reconstruction. Given the past and future reconstructed results $\hat{\mathbf{X}}_{rec}^{P}$ and $\hat{\mathbf{X}}_{rec}^{F}$, and the corresponding ground-truth (i.e. complete sequence) $\mathbf{X}_{rec}^{P}$ and $\mathbf{X}_{rec}^{F}$, the loss function can be represented as follow:
$\begin{split}\mathcal{L}_{pretrain}&=\frac{1}{TJ}\sum_{t=1}^{T}\sum_{j=1}^{J}||\mathbf{X}_{rec}^{P,(t,j)}-\hat{\mathbf{X}}_{rec}^{P,(t,j)}||_{F}^{2}\\ &+\frac{\alpha}{LJ}\sum_{l=1}^{L}\sum_{j=1}^{J}||\mathbf{X}_{rec}^{F,(l,j)}-\hat{\mathbf{X}}_{rec}^{F,(l,j)}||_{F}^{2}\end{split}$ (8)
@@
For finetune stage, given the predicted sequence $\hat{\mathbf{X}}$ and the ground-truth $\mathbf{X}$, the loss function can be defined as:
$\mathcal{L}_{finetune}=\frac{1}{LJ}\sum_{l=1}^{L}\sum_{j=1}^{J}||\mathbf{X}_{j}^{l}-\hat{\mathbf{X}}_{j}^{l}||_{F}^{2}$ (9)
@@
DMGNN[20] 17.3 30.7 54.6 65.2 11.0 21.4 36.2 43.9 9.0 17.6 32.1 40.3 17.3 34.8 61.0 69.8
MSR[7] 12.2 22.7 38.6 45.2 8.4 17.1 33.0 40.4 8.0 16.3 31.3 38.2 12.0 26.8 57.1 69.7
Mixer[2] 12.0 23.0 42.4 51.3 7.8 16.5 34.3 43.0 7.4 15.2 30.3 37.6 10.7 25.0 56.9 71.1
PGBIG[26] 10.2 19.8 34.5 40.3 7.0 15.1 30.6 38.1 6.6 14.1 28.2 34.7 10.0 23.8 53.6 66.7
SPGSN[24] 10.1 19.4 34.8 41.5 7.1 14.8 30.5 37.9 6.7 13.8 28.0 34.6 10.4 23.8 53.6 67.1
siMLPe[12] 11.9 19.3 31.9 36.9 10.2 16.5 30.7 37.3 10.4 16.4 29.7 35.9 12.7 25.3 55.1 68.1
DSTD-GC[10] 11.1 22.4 38.8 45.2 7.0 15.5 31.7 39.2 6.6 14.8 29.8 36.7 10.0 24.4 54.5 67.4
AuxFormer[41] 8.9 16.9 30.1 36.1 6.4 14.0 28.8 35.9 5.7 11.4 22.1 27.9 8.6 18.8 38.8 49.2
CIST-GCN[33] 11.8 23.4 40.5 46.5 6.7 14.8 29.8 36.8 7.3 15.6 31.0 38.0 10.2 23.7 52.3 65.3
@@
DMGNN[20] 13.1 24.6 64.7 81.9 23.3 50.3 107.3 132.1 12.5 25.8 48.1 58.3 15.3 29.3 71.5 96.7
MSR[7] 8.6 19.7 43.3 53.8 16.5 37.0 77.3 93.4 10.1 20.7 41.5 51.3 12.8 29.4 67.0 85.0
Mixer[2] 7.5 18.4 43.9 55.3 16.4 37.1 80.2 98.2 9.1 19.5 42.4 53.6 11.5 27.0 64.9 83.5
PGBIG[26] 7.2 17.6 40.9 51.5 15.2 34.1 71.6 87.1 8.3 18.3 38.7 48.4 10.7 25.7 60.0 76.6
SPGSN[24] 7.4 17.2 40.0 50.3 14.6 32.6 70.6 86.4 8.7 18.3 38.7 48.5 10.7 25.3 59.9 76.5
siMLPe[12] 11.0 19.4 41.6 52.1 16.2 33.7 70.9 86.3 11.9 20.3 39.7 48.9 13.9 27.3 62.0 78.6
DSTD-GC[10] 6.9 17.4 40.1 51.7 14.3 33.5 72.2 87.3 8.5 19.2 40.3 49.9 10.1 25.4 60.6 77.3
AuxFormer[41] 6.8 17.0 40.3 51.6 13.5 31.3 69.2 85.4 7.9 17.3 37.4 47.2 8.8 19.1 39.2 51.0
CIST-GCN[33] 7.3 18.1 43.6 55.3 13.7 31.0 65.7 79.9 8.6 18.5 39.3 49.6 9.6 23.7 57.7 75.0
@@
Motion Average Param(M)
Milliseconds 80 160 320 400 560 1000 -
DMGNN[20] 17.0 33.6 65.9 79.7 103.0 137.2 4.82
MSR[7] 12.1 25.6 51.6 62.9 81.1 114.2 6.30
Mixer[2] 11.1 24.0 51.5 64.0 83.5 116.5 0.03
PGBIG[26] 10.3 22.7 47.4 58.5 76.9 110.3 1.80
SPGSN[24] 10.4 22.3 47.1 58.3 77.1 108.8 5.66
siMLPe[12] 13.4 23.7 47.4 58.0 77.9 111.7 0.14
DSTD-GC[10] 10.4 23.3 48.8 59.8 77.8 111.0 0.18
AuxFormer[41] 9.5 20.6 43.4 54.1 75.3 107.0 1.00
CIST-GCN[33] 10.5 23.2 47.9 59.0 77.2 110.3 0.35
GCNext [39] 9.3 21.5 45.5 56.4 74.7 108.7 -
Ours 8.8 19.3 41.1 51.8 70.0 101.5 1.66
@@ 4.1 Dataset
Human3.6M[17] is a large-scale, accurately annotated 3D human pose dataset that captures the motions of 11 professional actors in daily life scenarios using motion capture technology, totaling over 3.6 million pose instances. Following previous works[26, 24], only 22 joints are used, focusing on those that exhibit significant movement, and all sequences are down-sampled to 25 fps.
@@
3DPW[37] stands out as a pioneering dataset in motion related tasks, primarily due to its focus on capturing human motion in natural settings outside of the typical laboratory environment. This dataset provides a rich collection of video sequences captured at a frame rate of 30 frames per second, which includes diverse activities such as ’Walking’, ’Running’, and interacting with various objects.
@@
AMASS[27] consolidates 15 existing motion capture datasets into a single, standardized framework. By converting these diverse datasets into a common representation, AMASS provides over 40 hours of meticulously annotated motion data. Following the previous works[34, 40], we discarded the detailed joints of the hand and sampled the sequence to 24fps.
Figure 3: Visualization results of short-term prediction (a) and reconstruction (b). The blue lines represent predicted/reconstructed motions, and the red lines represent ground-truths. We map the results to 2D space for display.
@@ 4.2 Experimental Settings
For the PME and the FMP, the number of modules (denoted as $N,M$) are both set to 3. The number of feature channels remains consistent at 128 throughout the network architecture. In the attention calculation, we employed 8 heads, with each head comprising 32 feature channels. During the pretraining phase, the loss weight (denoted as $\alpha$) is set to 1, indicating an equal contribution of loss from both past and future sequence reconstruction. We provide more implementation details in the supplementary materials (Sec. 8).
@@ 4.3 Comparison to State-of-the-art Methods
We compare our method with the recent state-of-the-art methods, including DMGNN (CVPR’20)[20, 30, 1, 7, 2, 26, 24, 12, 10, 41, 40, 33, 39]. To ensure fair comparisons, we re-evaluated some methods on the entire test set to align the testing standard using their official pretrained models, or retrain the model (Only for the cases where the official pre-training model is not provided) using their default hyperparameters on the Human3.6M.
@@
We use MPJPE (Mean Per Joint Position Error) as the evaluation metric, which is widely used in motion prediction tasks[26, 24, 41]. Tab. 1 shows the quantitative comparisons of short-term motion prediction on Human3.6M dataset, which aims to predict the poses within 400 milliseconds. Detailed actions are listed. As can be seen, our method outperforms the others on most actions, especially the ”Smoking”, ”Greeting”, and ”Posing”. Tab. 2 displays the average results of short - and long-term predictions for all actions on Human3.6M. Our approach effectively reduces prediction errors by 5.4%/10.2%/9.7%/8.2%/6.3%/6.6% at 80ms/160ms/320ms/400ms/560ms/1000ms.
Table 3: Comparisons of MPJPEs between our proposed method with other state-of-the-art methods for both short-term and long-term prediction on AMASS. The best results are highlighted in bold.
@@
LTD-10-10[28] 10.3 19.3 36.6 44.6 61.5 75.9 86.2 91.2
LTD-10-25[28] 11.0 20.7 37.8 45.3 57.2 65.7 71.3 75.2
HisRep[29] 11.3 20.7 35.7 42.0 51.7 58.6 63.4 67.2
Mao W, et al.[30] 11.0 20.3 35.0 41.2 50.7 57.4 61.9 65.8
siMLPe[12] 10.8 19.6 34.3 40.5 50.5 57.3 62.4 65.7
DD-GCN[40] 10.4 19.1 33.6 39.8 49.3 56.5 61.3 64.6
@@
Tab. 3 shows the comparisons on AMASS dataset. Our method achieves a reduction in prediction errors of 17.3%/19.9%/17.6%/16.6%/14.8%/14.5%/13.5%/13.3% at 80ms/160ms/320ms/400ms/560ms/720ms/880ms/1000ms. We found that our method exhibits the most significant improvement on this dataset. This can be attributed to the larger data size of AMASS compared to Human3.6M and 3DPW. The self-reconstruction pretraining enables our model to learn a more comprehensive representation of motion on such a large dataset.
@@
DMGNN[20] 17.80 37.11 70.38 94.12 109.67 123.93
HisRep[29] 15.88 35.14 66.82 93.55 107.63 114.75
MSR-GCN[7] 15.70 33.48 65.02 93.81 108.15 116.31
STSGCN[35] 18.32 37.79 67.51 92.75 106.65 112.22
PGBIG[26] 17.66 35.32 67.83 89.60 102.59 109.41
SPGSN[24] 15.39 32.91 64.54 91.62 103.98 111.05
AuxFormer[41] 14.21 30.04 58.50 89.45 100.78 107.45
@@
Table 5: Comparison of Motion Classification Using Network-Output Features
w/o Network AuxFormer[41] Ours
Accuracy 25.6% 80.8% 95.0%
@@
To validate the representation learning capability of our proposed method, we first extract motion features from the outputs of the trained model (outputs from the penultimate layer). Subsequently, we train a simple linear classifier to classify these features. As shown in Tab. 5, our method significantly improves classification accuracy compared to methods that do not involve feature extraction via neural networks and AuxFormer[41].
@@ 4.4 Visualization of Prediction and Reconstruction
Fig. 3 (a) presents the visualization of prediction results. We conducted comparisons with PGBIG[26] and Mixer[2] on challenging samples. In the ’Posing’ and ’Discussion’ actions, our model successfully captures significant upward/downward movements of the arms. This can be attributed to the utilization of past motion guidance, which enables the model to comprehend the influence of previous motion patterns on future sequences. We also provide reconstruction results with a mask rate of 75% in Fig. 3 (b). Similar to the property of mask-based pretraining[15], even with a high mask rate of 75% for joints, the model can still accurately reconstruct the exact location of them.

### v2

@@ 3.4 Fine-tuning for Prediction and Understanding
$\mathbf{H}^{F}$ $=\mathcal{F}_{JE}(\mathcal{F}_{FMD}(\mathbf{H}_{0}^{F},\mathbf{H}^{P}))$ (6)
$\mathbf{X}^{F}$ $=\mathcal{F}_{PH}(\mathbf{H}^{F}),$ (7)
where $\mathbf{H}_{0}^{F}$ is the zero-initialized future state and $\mathcal{F}_{JE}$ denotes the joint embedder.
Beyond motion prediction, our framework also supports future motion understanding. Given the past-context representation $\mathbf{H}^{P}$, we apply a Future Semantic Decoder (FSD) $\mathcal{F}_{FSD}$ to generate a future semantic description:
$\mathbf{Y}^{F}=\mathcal{F}_{FSD}(\mathbf{H}^{F},\mathbf{H}^{P}),$ (8)
where $\mathbf{Y}^{F}=[y_{1},\ldots,y_{M}]$ is the predicted future text sequence. This enables the model to jointly perform future motion prediction and future motion understanding within a unified fine-tuning pipeline.
@@ 3.5 Loss Function
Loss Function of Pretraining Stage
Pretraining stage performs past sequence reconstruction and future sequence reconstruction. Given the past and future reconstructed results $\hat{\mathbf{X}}_{rec}^{P}$ and $\hat{\mathbf{X}}_{rec}^{F}$, and the corresponding ground-truth (i.e. complete sequence) $\mathbf{X}_{rec}^{P}$ and $\mathbf{X}_{rec}^{F}$, the loss function can be represented as follow:
[…]
@@
In the finetuning stage, the overall objective integrates motion reconstruction and semantic prediction into a single loss:
$\mathcal{L}_{finetune}$ $=\underbrace{\frac{1}{LJ}\sum_{l=1}^{L}\sum_{j=1}^{J}\left\|\mathbf{X}_{j}^{l}-\hat{\mathbf{X}}_{j}^{l}\right\|_{F}^{2}}_{\mathcal{L}_{motion}}$ (10)
$\quad+\lambda\underbrace{\left(-\sum_{d=1}^{D}\log p(\mathbf{Y}^{F}_{d}\mid\hat{\mathbf{Y}}^{F})\right)}_{\mathcal{L}_{text}}.$
Jointly optimizing the two losses encourages mutually reinforcing representations, leading to more accurate future motion prediction and more consistent semantic generation.
@@
DSTD-GC [fu2023learning] 11.1 22.4 38.8 45.2 7.0 15.5 31.7 39.2 6.6 14.8 29.8 36.7 10.0 24.4 54.5 67.4
AuxFormer [xu2023auxiliary] 8.9 16.9 30.1 36.1 6.4 14.0 28.8 35.9 5.7 11.4 22.1 27.9 8.6 18.8 38.8 49.2
CIST-GCN [medina2024context] 11.8 23.4 40.5 46.5 6.7 14.8 29.8 36.8 7.3 15.6 31.0 38.0 10.2 23.7 52.3 65.3
@@
DSTD-GC [fu2023learning] 6.9 17.4 40.1 51.7 14.3 33.5 72.2 87.3 8.5 19.2 40.3 49.9 10.1 25.4 60.6 77.3
AuxFormer [xu2023auxiliary] 6.8 17.0 40.3 51.6 13.5 31.3 69.2 85.4 7.9 17.3 37.4 47.2 8.8 19.1 39.2 51.0
CIST-GCN [medina2024context] 7.3 18.1 43.6 55.3 13.7 31.0 65.7 79.9 8.6 18.5 39.3 49.6 9.6 23.7 57.7 75.0
@@
Motion Average
Milliseconds 80 160 320 400 560 1000
DMGNN [li2020dynamic] 17.0 33.6 65.9 79.7 103.0 137.2
MSR [dang2021msr] 12.1 25.6 51.6 62.9 81.1 114.2
Mixer [bouazizi2022motionmixer] 11.1 24.0 51.5 64.0 83.5 116.5
PGBIG [ma2022progressively] 10.3 22.7 47.4 58.5 76.9 110.3
SPGSN [li2022skeleton] 10.4 22.3 47.1 58.3 77.1 108.8
siMLPe [guo2023back] 13.4 23.7 47.4 58.0 77.9 111.7
DSTD-GC [fu2023learning] 10.4 23.3 48.8 59.8 77.8 111.0
AuxFormer [xu2023auxiliary] 9.5 20.6 43.4 54.1 75.3 107.0
CIST-GCN [medina2024context] 10.5 23.2 47.9 59.0 77.2 110.3
GCNext [wang2024gcnext] 9.3 21.5 45.5 56.4 74.7 108.7
ALIEN [wei2025alien] 9.7 21.8 47.0 58.3 - -
Ours 8.8 19.3 41.1 51.8 70.0 101.5
@@ 4.1 Dataset
Human3.6M [ionescu2013human3] is a large-scale, accurately annotated 3D human pose dataset that captures the motions of 11 professional actors in daily life scenarios using motion capture technology, totaling over 3.6 million pose instances. Following previous works [ma2022progressively, li2022skeleton], only 22 joints are used, focusing on those that exhibit significant movement, and all sequences are down-sampled to 25 fps.
@@
3DPW [von2018recovering] stands out as a pioneering dataset in motion related tasks, primarily due to its focus on capturing human motion in natural settings outside of the typical laboratory environment. This dataset provides a rich collection of video sequences captured at a frame rate of 30 frames per second, which includes diverse activities such as ‘Walking’, ‘Running’, and interacting with various objects.
@@
AMASS [mahmood2019amass] consolidates 15 existing motion capture datasets into a single, standardized framework. By converting these diverse datasets into a common representation, AMASS provides over 40 hours of meticulously annotated motion data. Following the previous works [shi2023multi, wang2024dynamic], we discarded the detailed joints of the hand and sampled the sequence to 24fps.
FineMotion
To incorporate fine-grained action semantics, we adopt FineMotion [^1] https://github.com/CVI-SZU/FineMotion, an open-source extension of HumanML3D that re-labels all motions with detailed body-part movement descriptions. FineMotion segments each sequence into 0.5-second snippets and provides both human-annotated and automatically generated descriptions. Based on this resource, we construct our prediction dataset using a 2-second sliding window (0.5s past motion and 1.5s future motion). For each window, we concatenate the snippet-level descriptions corresponding to the future segment to form its semantic target. This process yields 12,628 human-annotated samples and 248,126 automatically annotated samples, which we split into train/val/test sets following an 80% / 5% / 15% ratio.
Figure 4: Visualization results of motion prediction. The red lines represent predicted motions, and the blue lines represent ground-truths.
Figure 5: Visualization results of motion reconstruction on H36M. The blue lines represent reconstructed motions, and the red lines represent ground-truths.
Figure 6: Visualization of motion prediction and understanding.
@@ 4.2 Experimental Settings
For the PME, FMD and FSU, the number of modules are all set to 3. The number of feature channels remains consistent at 128 throughout the PME and FMD, and remains at 256 throughout the FSU. In the attention calculation, we employed 8 heads, with each head comprising 32 feature channels. During the pretraining phase, the loss weight (denoted as $\alpha$) is set to 1, indicating an equal contribution of loss from both past and future sequence reconstruction. During the finetune phase, the loss weight (denoted as $\lambda$) is set to 0.1. The parameters in FSU are updated every 5 epoch.
@@ 4.3 Comparison to State-of-the-art Methods
We compare our method with the recent state-of-the-art methods [li2020dynamic, mao2021multi, aksan2021spatio, dang2021msr, bouazizi2022motionmixer, ma2022progressively, li2022skeleton, guo2023back, fu2023learning, xu2023auxiliary, wang2024dynamic, medina2024context, wang2024gcnext, wei2025alien]. To ensure fair comparisons, we re-evaluated some methods on the entire test set to align the testing standard using their official pretrained models, or retrain the model (Only for the cases where the official pre-training model is not provided) using their default hyperparameters on the Human3.6M.
@@
We use MPJPE (Mean Per Joint Position Error) as the evaluation metric, which is widely used in motion prediction tasks [ma2022progressively, li2022skeleton, xu2023auxiliary]. Tab. 1 shows the quantitative comparisons of short-term motion prediction on Human3.6M dataset, which aims to predict the poses within 400 milliseconds. Detailed actions are listed. As can be seen, our method outperforms the others on most actions, especially the “Smoking”, “Greeting”, and “Posing”. Tab. 2 displays the average results of short- and long-term predictions for all actions on Human3.6M. Our approach effectively reduces prediction errors by 5.4%/10.2%/9.7%/8.2%/6.3%/6.6% at 80ms/160ms/320ms/400ms/560ms/1000ms.
Table 3: Comparisons of MPJPEs between our proposed method with other SOTA methods for both short-term and long-term prediction on AMASS. The best results are highlighted in bold.
@@
LTD-10-10 [mao2019learning] 10.3 19.3 36.6 44.6 61.5 75.9 86.2 91.2
LTD-10-25 [mao2019learning] 11.0 20.7 37.8 45.3 57.2 65.7 71.3 75.2
HisRep [mao2020history] 11.3 20.7 35.7 42.0 51.7 58.6 63.4 67.2
Mao W, et al. [mao2021multi] 11.0 20.3 35.0 41.2 50.7 57.4 61.9 65.8
siMLPe [guo2023back] 10.8 19.6 34.3 40.5 50.5 57.3 62.4 65.7
DD-GCN [wang2024dynamic] 10.4 19.1 33.6 39.8 49.3 56.5 61.3 64.6
@@
Tab. 3 shows the comparisons on AMASS dataset. Our method achieves a reduction in prediction errors of 15.9%. We found that our method exhibits the most significant improvement on this dataset. This can be attributed to the larger data size of AMASS compared to Human3.6M and 3DPW. The self-reconstruction pretraining enables our model to learn a more comprehensive representation of motion on such a large dataset.
@@
DMGNN [li2020dynamic] 17.80 37.11 70.38 94.12 109.67 123.93
HisRep [mao2020history] 15.88 35.14 66.82 93.55 107.63 114.75
MSR-GCN [dang2021msr] 15.70 33.48 65.02 93.81 108.15 116.31
PGBIG [ma2022progressively] 17.66 35.32 67.83 89.60 102.59 109.41
SPGSN [li2022skeleton] 15.39 32.91 64.54 91.62 103.98 111.05
AuxFormer [xu2023auxiliary] 14.21 30.04 58.50 89.45 100.78 107.45
ALIEN [wei2025alien] 14.83 32.59 62.06 90.25 103.62 -
@@
Comparisons on Joint Prediction and Understading
As shown in Tab. 6, our method consistently achieves lower MPJPEs on the processed FineMotion dataset across all prediction horizons compared to prior approaches. Notably, introducing the understanding branch further improves motion prediction quality, indicating that capturing semantic cues of future movements can, in turn, benefit joint-level forecasting.
As shown in Tab. 7, the IDs “A” and “B” refer to models trained on auto-labeled and human-labeled text annotations, respectively. Our lightweight semantic prediction head achieves competitive or superior performance across several key metrics when compared with MGMotionLLM [wu2025mg], despite the latter being a large-scale language model designed for motion–text generation. Moreover, when the motion prediction loss is removed, the performance on future text generation also degrades, indicating that accurate motion forecasting provides essential guidance for understanding future semantic.
Table 5: Comparison of classification accuracy and clustering quality on Human3.6M using network-output features. CLS accuracy evaluates discriminativeness, while inter- and intra-class distances measure feature separability.
Task Metrics w/o Network AuxFormer [xu2023auxiliary] Ours
[…]
@@
We evaluate the quality of the learned motion representation by extracting penultimate-layer features and training a lightweight linear classifier on them. As shown in Table 5, our method achieves substantially higher classification accuracy than both w/o Network and AuxFormer [xu2023auxiliary], demonstrating stronger feature discriminativeness. In addition, our features present larger inter-class distances and smaller intra-class distances, indicating better clustering behavior and clearer semantic separability. For the clustering metrics, we randomly sample 100 instances from each of the 15 action categories in Human3.6M to compute the intra- and inter-class distances.
@@ 4.4 Visualization of Prediction and Reconstruction
Fig. 4 presents the visualization of prediction results. We conducted comparisons with GCNext [wang2024gcnext] on challenging samples. Our model successfully captures significant movements of the arms and legs. This can be attributed to the utilization of past motion guidance, which enables the model to comprehend the influence of previous motion patterns on future sequences. Fig. 6 showcases the joint results of future motion prediction and future motion understanding, where our method consistently outputs motion sequences and semantic descriptions that are highly aligned. Furthermore, We provide reconstruction results with a mask rate of 75% in Fig. 5. Similar to the property of mask-based pretraining [he2022masked], even with a high mask rate of 75% for joints, the model can still accurately reconstruct the location of them.
Table 6: Comparisons of MPJPEs between our proposed method with other state-of-the-art methods on the processed FineMotion dataset. The best results are highlighted in bold.
Average MPJPE
Milliseconds 300 600 900 1200 1500
AuxFormer [xu2023auxiliary] 36.07 62.31 76.95 86.11 94.22
GCNext [wang2024gcnext] 44.05 73.19 88.14 97.03 103.9
Ours w/o Und. 34.96 57.39 69.08 76.58 84.95
Ours 34.69 56.71 67.99 75.33 83.65
Table 7: Comparisons of future motion understanding between our proposed method with MGMotionLLM [wu2025mg] on the processed FineMotion dataset. The best results are highlighted in bold.
ID Method BLEU@4 BLEU@1 ROUGE CIDEr
A MGMotionLLM [wu2025mg] 25.3 15.8 38.0 12.5
Ours w/o Prediction 15.6 7.9 31.2 36.2
Ours 17.1 9.4 33.5 47.6
B MGMotionLLM [wu2025mg] 13.1 7.5 35.8 3.3
Ours w/o Prediction 15.0 7.1 31.7 25.8
Ours 18.2 7.8 32.5 24.2